OpenStudy (aminul_ahasun):

# Find the slope of the tangent line to the curve at the given points in two ways: first by solving for y in terms of x and differentiating and then by implicit differentiation. * x^2 + y^2 = 1; (1/2,√3/2), (1/2,−√3/2)

OpenStudy (welshfella):

have you done any work on this so far?

OpenStudy (anonymous):

x^2 + y^2 = 1 y²=1-x^² y=\[\sqrt{1-x²}\] y=(1-x²)\[^{1/2}\] implicit method: d/dx (x²+y²=1) 2x+2yy'=0 y'=-2x/2y=-x/y =1/2\[(1-x²)^{-1/2}\](-2x) =\[-x/\sqrt{1-x²}\]

OpenStudy (welshfella):

to find the values of the slope plug the values of x into the derivative -x / sqrt(1 - x^2)